# Electricity Market Bidding for Renewable Electrolyzer Plants: An Opportunity Cost Approach

### v2

@@ III-A Case Study Description
III-B Analysis of the Merit Order Curve
@@
III-C Simulation and Benchmarking Framework
To analyze the impact of a market bidding REP on the system and benchmark its effects against a more traditional style of operation, we simulate the electricity market clearing by solving the DC OPF problem (9) for each hour of the year. To solve the DC OPF, we use the implementation provided in the Julia package PowerModels.jl [20]. Our source codes are publicly available in [21].
III-C1 REP Representations
@@
Market bidding REP: The REP participates in the market, submitting its opportunity cost curve. A DC OPF is solved for the RTS-GMLC with the cost curve of wind farm 156 altered to represent the presence of the electrolyzer as described in Section II-C. This case represents how the REP would behave when participating in the electricity market and fully exposed to the LMPs of the system (i.e., fully flexible).
Fixed consumption REP: We compare the market bidding REP to an REP where the electrolyzer is consuming at a fixed, constant power. To create a fair comparison, we define the fixed consumption power as the average consumption of the market bidding REP, i.e., the total electricity consumption of the market bidding REP over the year, divided by the number of hours in the year. The electrolyzer is modeled as an additional load at the REP’s node, not modifying the bid curve of wind farm 156. This new load is set to the calculated fixed consumption and is inflexible. However, for several hours, this level of consumption causes the DC OPF to become infeasible. In those hours, we first model the fixed consumption as a bid with a high marginal cost (of \$10,000 per MWh), thus approximating the highest electrolyzer consumption possible while maintaining a feasible program. We then rerun the DC OPF with the new load set equal to this (lower) consumption.
III-D System-focused Results
Figure 5: Total cost of generation and cost of generation per load for a fixed consumption and market bidding electrolyzer under the hydrogen prices of \$currency-dollar\\$\$1.5 and \$currency-dollar\\$\$6 per kg.
@@ III-D1 Cost of generation
We observe that the total cost of generation (two left plots) increases as the electrolyzer capacity increases. This is expected as the electrolyzer imposes an additional load on the system. For the smaller electrolyzer capacity of 100 MW, the cost increases by less than 2.5% for any operational mode or hydrogen price. However, for larger electrolyzer capacities, the total cost of generation increases by 10-20%. In general, the market bidding REP yields a smaller increase in the total cost of generation compared to the corresponding fixed consumption case. This is expected as the market bidding REP will reduce its consumption if the LMP is sufficiently high. However, for an electrolyzer capacity of 1000 MW operating at the higher hydrogen price, the market bidding REP causes a higher cost of generation than the fixed consumption counterpart. This is because the fixed consumption REP is consuming significantly less power overall due to congestion in the network, as will be elaborated later in Section III-E1.
Next, we consider the cost of generation per load, shown in the two plots to the right in Fig. 5. We first observe that adding a small electrolyzer with 100 MW capacity decreases the cost of generation per load for both hydrogen prices and REP representations. This is because part of added REP consumption is met by otherwise curtailed renewable power. For electrolyzer capacities of 500 and 1000 MW, the average cost generally increases. However, the magnitude of the increase depends on the case. Cases with fixed consumption and/or a higher hydrogen price lead to significant increases in average cost, reaching up to 7. 7% (although a market bidding REP causes smaller increases compared to a fixed consumption REP). In contrast, adding a market bidding REP with a lower hydrogen price yields a negligible increase in generation cost per load.
From these results, we conclude that to keep the cost of generation low, the REP must be sensitive to the LMP, i.e., operated as a market bidder. Further, higher hydrogen prices reduce the benefits of market bidding REP operation. Following the merit order curve in Fig. 4, the market bidding REP with a high marginal cost is only dispatched if there is network congestion, and will therefore impact the total cost of generation in a way similar to that of a fixed consumption REP. We observe this effect in the cases with a hydrogen price of \$6 per kg and large electrolyzer capacities, and note that subsidies for hydrogen production, through inflating the hydrogen price, might result in less flexible operation of the REP and thus increase the cost of generation of its connected system.
@@
III-D2 Renewable power curtailment
We next investigate how the presence of an electrolyzer in the system affects the curtailment of renewable power, as it is expected that electrolyzers can use surplus renewable power and thus reduce the need for curtailment. Fig. 6 shows the total renewable power curtailment over the year in the entire system, for the different electrolyzer capacities, REP representations, and hydrogen prices explored previously. Intuitively, as the load in the system increases, the amount of curtailed renewable power should decrease. This trend is observed in the market bidding cases with a low hydrogen price. However, in the other cases, increasing the electrolyzer capacity from 500 to 1000 MW actually increases the curtailment. This increase in curtailment is a result of congestion in the network. When the REP bids to import power at a sufficiently high price (or in the fixed consumption cases), congestion forces the DC OPF to dispatch expensive, non-renewable generation and curtails renewable generation to allow the REP to import more power.
We draw a conclusion similar to that in Section III-D1, that high hydrogen prices reduce the flexibility of electrolyzers, potentially leading to negative impacts on RES curtailment.
III-D3 System emissions
@@
Our results show that for a given electrolyzer capacity, both the total emissions and emissions per load remain similar irrespective of both how the REP is operated and the hydrogen price. This result is understood by examining the merit order curve of the system in Fig. 4. The cheapest generation sources, i.e., renewable and nuclear capacities, are indeed also the least emission intense. However, the (generally) cheapest conventional generation capacities, the coal-fueled plants, have higher emission factors than the more expensive natural gas and oil-fueled plants. Therefore, following price-signals cannot guarantee lower system emissions.
@@ III-E1 Electrolyzer utilization
For an electrolyzer with 1000 MW capacity, we observe that the average load is significantly below capacity for the market bidding REPs, even for the case with a high hydrogen price. This indicates that a combination of congestion and high prices at the REP node reduces the viability of importing sufficient power to run the electrolyzer at full capacity.
@@
Further, while the fixed consumption REP should consume the average electrolyzer load of the corresponding market bidding case, we observe that the actual average consumption is significantly lower than for the market bidding REP. As previously explained, the fixed consumption level leads to an infeasible program in certain hours, and for those hours we reduce the consumption level of the fixed electrolyzer. The lower average load therefore indicates that a fixed consumption at this level is frequently infeasible.
III-E2 Profit
@@
We next discuss the REP’s operational profit under the different cases, shown in the second to last column of Table II. Recall that the REP profit is calculated assuming that the REP imports and exports power at a price equal to the LMP for all cases. Under this assumption, we hypothesize that the market bidding REP will achieve a profit equal to or higher than that of the fixed consumption REP, since the market bidding REP is actively adapted to the LMP. From our results, we observe that this hypothesis holds true. For the smaller electrolyzer capacity and high hydrogen price, the profit is similar for market bidding and fixed consumption operation, as it is cost-effective to run the electrolyzer at full capacity in both cases. However, if the hydrogen price is lower and/or the electrolyzer capacity is sufficiently large (e.g., 500 or 1000 MW) to cause congestion or infeasibility, we observe that there can be significant differences in profit between the market bidding and fixed consumption cases. In the case of a 500 MW electrolyzer capacity and a lower hydrogen price, the market bidding case yields a 26% profit increase compared to the fixed consumption case. In the case of a 1000 MW electrolyzer capacity, the profits achieved through active market bidding far exceed those of a fixed consumption, which both exhibit a loss (i.e., a negative profit).
@@
Finally, we observe that a larger electrolyzer is not always better in terms of profit. When we compare profits for market bidding REPs with 500 and 1000 MW electrolyzer capacity, we observe that a larger capacity increases profits in the case with a \$1.5 per kg hydrogen price but decreases profits in the case with a \$6 per kg hydrogen price. The average LMP at the REP’s node, given by the last column of Table II, shows that there is a significant increase in LMP when the electrolyzer capacity increases from 500 to 1000 MW in certain cases. We examine the case of a market bidding REP under the higher hydrogen price. In this case, the LMP increases from \$31.77 to \$81.54 per MWh. Recalling the merit order curve of Fig. 4, the marginal cost of the REP under a higher hydrogen price lies in the range of \$80-120 per MWh. The REP therefore makes more profit from producing hydrogen than from selling electricity at LMPs up to this range. However, with increasing LMP, the profit associated with hydrogen production decreases. The increase in electrolyzer capacity when operating under the higher hydrogen price increases the LMP to a point where the REPs profit decreases.
@@
We conclude that to evaluate the profit of a market bidding REP, it could be essential to model the on the price formation process through, e.g., a DC OPF, as the presence of the electrolyzer in the system can highly impact the LMPs and the REP’s profit. This becomes more important with larger electrolyzer capacities and higher hydrogen prices.
@@
III-F The Effect of Transmission Network Modeling
Finally, we investigate the effect of transmission network modeling on the system- and REP-focused results. We compare three network models with varying levels of detail regarding transmission network constraints: a nodal model (which includes all transmission lines), a zonal model (which relaxes all intra-regional transmission constraints), and a copper-plate model (which relaxes all transmission constraints). To create the zonal model, we define each of the three regions in the RTS-GMLC system, as described by Fig. 3, as individual zones. The total inter-regional transmission capacities between regions 1$\leftrightarrow$2, 1$\leftrightarrow$3, and 2$\leftrightarrow$3, are 1175, 500, and 500 MW, respectively. We solve the DC OPF problem for the system with (i) no REP (base case) and (ii) a market bidding, 1000 MW electrolyzer REP with low hydrogen price, and (iii) a market bidding, 1000 MW electrolyzer REP with high hydrogen price. These three cases are solved for each transmission network model, resulting in a total of nine cases. For each case, we compute the system load, cost of generation per load, system emissions per load, and total renewable power curtailment. The system-focused results are reported in Table III. We further report REP-focused results including the average electrolyzer load, REP’s profit, and the average LMP at the REP’s node in Table IV. All results are computed in the same manner as described in Section III-D.
III-F1 System-focused results
We first examine the system load (first column of Table III). For both hydrogen prices, the increase in system load from the base system is lower under a nodal model than under a zonal or copper-plate model. In the case of the higher hydrogen price, the consumption of the REP increases the system load by 23% under a zonal model and by 17% under a nodal model. This is expected, as our analysis in Section III-E showed that the 1000 MW electrolyzer utilization is limited by system congestion and increased LMPs under the nodal model. However, there is no substantial further increase in system load when going from a zonal model to a copper-plate model at any hydrogen price. This suggests that congestion is primarily intra-regional and can only be fully captured by a nodal model.
@@ III-F1 System-focused results
We conclude that the choice of transmission network model significantly affects the perceived results of adding an electrolyzer to a power system, as the added load could induce congestion in the system.
@@
III-F2 REP-focused results
The results for electrolyzer utilization (first column of Table IV) further illustrate to what extent the relaxed grid models overestimate the power consumption of the REP. For example, in the case with a higher hydrogen price, the electrolyzer utilization is at 100% under the relaxed network models, compared to 73% under the nodal model. This increase in utilization contributes to an overestimation of the REP’s profits, which are reported in the second column. Further, the increase in perceived profit is in part due to the LMPs at the REP’s node (last column of Table III). The copper-plate and zonal models yield lower average LMPs, as reduced congestion allows for the dispatch of cheaper generators. In the case of the higher hydrogen price, this leads to vastly differing results on REP profitability. Specifically, the zonal and copper-plate models result in more than twice the REP’s profit compared to the nodal model. The results presented for a nodal model in Section III-E show that a 1000 MW electrolyzer would actually yield a lower profit than a 500 MW electrolyzer in the case of a high hydrogen price, due to the increase in LMPs. This could dis-incentivize larger electrolyzer capacity investments under a nodal pricing scheme. However, the results of profits in Table IV show that this incentive is lost under the zonal and copper-plate models.
@@ III-F2 REP-focused results
IV Conclusions
This paper proposes an opportunity-cost curve for REPs, with the aim of facilitating the integration of such plants in the electricity market. The opportunity cost is derived from the revenue of hydrogen production and forms a curve that can be used to model REP bids in an electricity market. We compare a market bidding REP, bidding its derived opportunity-cost curve, to a REP with a fixed electrolyzer consumption. The results are obtained by solving a DC OPF of the RTS-GMLC test system, modifying a pre-existing wind farm cost curve to represent the REP.
The results show that active market participation can efficiently mitigate increases in the cost of generation per load that occur when the electrolyzer is operated in a fixed manner. Similarly, a market bidding REP efficiently reduces the curtailed renewable power in the system. However, high hydrogen prices will move the REP further up on the merit-order curve, causing it to behave more like a fixed load in the system. Further, market bidding in itself is not necessarily an efficient measure to prevent increases in system emissions due to the lack of correlation between the emission intensities and the marginal cost of generators.

### v3

@@ IV-A Case Study Description
IV-B Analysis of the Merit Order Curve
@@
IV-C Simulation and Benchmarking Framework
To analyze the impact of a market-bidding REP on the system and benchmark its effects against a more traditional style of operation, we simulate the electricity market clearing by solving the DC OPF problem (13) for each hour of the year. To solve the DC OPF, we use the implementation provided in the Julia package PowerModels.jl [38]. Our source codes are publicly available in [39].
IV-C1 REP Representations
@@
Market-bidding REP: The REP participates in the market, submitting its opportunity cost curve. A DC OPF is solved for the RTS-GMLC with the cost curve of wind farm 156 altered to represent the presence of the electrolyzer as described in Section III-D. This case represents how the REP would behave when participating in the electricity market and fully exposed to the LMPs of the system (i.e., fully flexible).
Fixed consumption REP: We compare the market-bidding REP to an REP where the electrolyzer is consuming at a fixed, constant power. To create a fair comparison, we define the fixed consumption power as the average consumption of the market-bidding REP, i.e., the total electricity consumption of the market-bidding REP over the year, divided by the number of hours in the year. The electrolyzer is modeled as an additional load at the REP’s node, not modifying the bid curve of wind farm 156. This new load is set to the calculated fixed consumption and is inflexible. However, for several hours, this level of consumption causes the DC OPF to become infeasible. In those hours, we first model the fixed consumption as a bid with a high marginal cost (of \$10,000 per MWh), thus approximating the highest electrolyzer consumption possible while maintaining a feasible program. We then rerun the DC OPF with the new load set equal to this (lower) consumption.
IV-D System-focused Results
Figure 7: Total cost of generation and cost of generation per load for a fixed consumption and market-bidding electrolyzer under the hydrogen prices of \$currency-dollar\\$\$1.5 and \$currency-dollar\\$\$6 per kg.
@@ IV-D1 Cost of generation
We observe that the total cost of generation (two left plots) increases as the electrolyzer capacity increases. This is expected as the electrolyzer imposes an additional load on the system. For the smaller electrolyzer capacity of 100 MW, the cost increases by less than 2.5% for any operational mode or hydrogen price. However, for larger electrolyzer capacities, the total cost of generation increases by 10-20%. In general, the market-bidding REP yields a smaller increase in the total cost of generation compared to the corresponding fixed consumption case. This is expected as the market-bidding REP will reduce its consumption if the LMP is sufficiently high.
Next, we consider the cost of generation per load, shown in the two plots to the right in Fig. 7. We first observe that adding a small electrolyzer with 100 MW capacity decreases the cost of generation per load for both hydrogen prices and REP representations. This is because part of added REP consumption is met by otherwise curtailed renewable power. For electrolyzer capacities of 500 and 1000 MW, the average cost generally increases. However, the magnitude of the increase depends on the case. Cases with fixed consumption and/or a higher hydrogen price lead to significant increases in average cost, reaching up to 7.7% (although a market-bidding REP causes smaller increases compared to a fixed consumption REP). In contrast, adding a market-bidding REP with a lower hydrogen price yields a negligible increase in generation cost per load.
From these results, we conclude that to keep the cost of generation low, the REP must be sensitive to the LMP, i.e., operated as a market bidder. Further, higher hydrogen prices reduce the benefits of market-bidding REP operation. Following the merit order curve in Fig. 6, the market-bidding REP with a high marginal cost is only dispatched if there is network congestion, and will therefore impact the total cost of generation in a way similar to that of a fixed consumption REP. We observe this effect in the cases with a hydrogen price of \$6 per kg and large electrolyzer capacities, and note that subsidies for hydrogen production, through inflating the hydrogen price, might result in less flexible operation of the REP and thus increase the cost of generation of its connected system.
@@
IV-D2 Renewable power curtailment
We next investigate how the presence of an electrolyzer in the system affects the curtailment of renewable power, as it is expected that electrolyzers can use surplus renewable power and thus reduce the need for curtailment. Fig. 8 shows the total renewable power curtailment over the year in the entire system, for the different electrolyzer capacities, REP representations, and hydrogen prices explored previously. In absence of congestion, increasing system load should yield decreasing renewable curtailment due to the renewable generators’ position on the merit order curve of Fig. 6. This trend is observed in the market-bidding cases with a low hydrogen price. However, in the other cases, increasing the electrolyzer capacity from 500 to 1000 MW actually increases the curtailment. This increase in curtailment is a result of congestion in the network. When the REP bids to import power at a sufficiently high price (or in the fixed consumption cases), congestion forces the DC OPF to dispatch expensive, non-renewable generation and curtails renewable generation to allow the REP to import more power.
We draw a conclusion similar to that in Section IV-D1, that high hydrogen prices reduce the flexibility of electrolyzers, potentially leading to negative impacts on RES curtailment.
IV-D3 System emissions
@@
Our results show that for a given electrolyzer capacity, both the total emissions and emissions per load remain similar irrespective of both how the REP is operated and the hydrogen price. Hence, there is no clear difference in the additional emissions from the system depending on the operational strategy of the REP, but rather depending on its overall consumption. This result is understood by examining the merit order curve of the system in Fig. 6. The cheapest generation sources, i.e., renewable and nuclear capacities, are indeed also the least emission intense. However, the (generally) cheapest conventional generation capacities, the coal-fueled plants, have higher emission factors than the more expensive natural gas and oil-fueled plants. Therefore, the market-bidding REP, while following price-signals from the electricity market, does not guarantee lower system emissions compared to the fixed REP.
@@ IV-E1 Electrolyzer utilization
For an electrolyzer with 1000 MW capacity, we observe that the average load is significantly below capacity for the market-bidding REPs, even for the case with a high hydrogen price. This indicates that a combination of congestion and high prices at the REP node reduces the viability of importing sufficient power to run the electrolyzer at full capacity.
@@
Further, while the fixed consumption REP should consume the average electrolyzer load of the corresponding market-bidding case, we observe that the actual average consumption is significantly lower than for the market-bidding REP. As previously explained, the fixed consumption level leads to an infeasible program in certain hours, and for those hours we reduce the consumption level of the fixed electrolyzer. The lower average load therefore indicates that a fixed consumption at this level is frequently infeasible.
IV-E2 Profit
@@
We next discuss the REP’s operational profit under the different cases, shown in the second to last column of Table II. Recall that the REP profit is calculated assuming that the REP imports and exports power at a price equal to the LMP for all cases. Under this assumption, we hypothesize that the market-bidding REP will achieve a profit equal to or higher than that of the fixed consumption REP, since the market-bidding REP is actively adapted to the LMP. From our results, we observe that this hypothesis holds true. For the smaller electrolyzer capacity and high hydrogen price, the profit is similar for market-bidding and fixed consumption operation, as it is cost-effective to run the electrolyzer at full capacity in both cases. However, if the hydrogen price is lower and/or the electrolyzer capacity is sufficiently large (e.g., 500 or 1000 MW) to cause congestion or infeasibility, we observe that there can be significant differences in profit between the market-bidding and fixed consumption cases. In the case of a 500 MW electrolyzer capacity and a lower hydrogen price, the market-bidding case yields a 26% profit increase compared to the fixed consumption case. In the case of a 1000 MW electrolyzer capacity, the profits achieved through active market bidding far exceed those of a fixed consumption, which both exhibit a loss (i.e., a negative profit).
While the market-bidding REP guarantees non-negative operational profits, it must still cover its investment costs over the long term. Thus, long-term profitability depends on operational profits being sufficient to recover the costs of constructing the wind farm and electrolyzer over the plant’s lifetime. Since this study focuses on the short term and does not consider factors such as optimal sizing and placement of REP components, we do not analyze long-term profitability. However, our results indicate a tendency for operational profits to decrease as electrolyzer capacity increases, which may serve as an important signal for investment decisions.
@@
When we compare profits for market-bidding REPs with 500 and 1000 MW electrolyzer capacity, we observe that a larger capacity increases profits in the case with a \$1.5 per kg hydrogen price but decreases profits in the case with a \$6 per kg hydrogen price. The average LMP at the REP’s node, given by the last column of Table II, shows that there is a significant increase in LMP when the electrolyzer capacity increases from 500 to 1000 MW in certain cases. We examine the case of a market-bidding REP under the higher hydrogen price. In this case, the LMP increases from \$31.77 to \$81.54 per MWh. Recalling the merit order curve of Fig. 6, the marginal cost of the REP under a higher hydrogen price lies in the range of \$80-120 per MWh. The REP therefore makes more profit from producing hydrogen than from selling electricity at LMPs up to this range. However, with increasing LMP, the profit associated with hydrogen production decreases. The increase in electrolyzer capacity when operating under the higher hydrogen price increases the LMP to a point where the REPs profit decreases.
Typically, a generator’s marginal cost curve reflects the electricity price at which it breaks even for a given output. Hence, if the cleared price is above its marginal cost, it makes a profit. Essential to the REP is that when wind production is available, the marginal opportunity cost curve does not represent where the profits are zero. Instead, it represents the marginal profit that the REP can draw from producing hydrogen. Whenever the REP has wind production available, it is guaranteed a minimum profit from hydrogen production. The REP can increase this profit through selling to the electricity market at high prices, or through buying additional electricity for hydrogen production at low prices. Therefore, an REP with wind production available, will generally draw additional profits through market interaction when the electricity price is different from its marginal cost curve. If wind production is not available, it only draws profit from prices lower than its marginal cost.
@@
We conclude that to evaluate the profit of a market-bidding REP, it could be essential to model the on the price formation process through, e.g., a DC OPF, as the presence of the electrolyzer in the system can highly impact the LMPs and the REP’s profit. This becomes more important with larger electrolyzer capacities and higher hydrogen prices.
@@
IV-F The Effect of Transmission Network Modeling
Finally, we investigate the effect of transmission network modeling on the system- and REP-focused results. We compare three network models with varying levels of detail regarding transmission network constraints: a nodal model (which includes all transmission lines), a zonal model (which relaxes all intra-regional transmission constraints), and a copper-plate model (which relaxes all transmission constraints). To create the zonal model, we define each of the three regions in the RTS-GMLC system, as described by Fig. 5, as individual zones. The total inter-regional transmission capacities between regions 1$\leftrightarrow$2, 1$\leftrightarrow$3, and 2$\leftrightarrow$3, are 1175, 500, and 500 MW, respectively. We solve the DC OPF problem for the system with (i) no REP (base case) and (ii) a market-bidding, 1000 MW electrolyzer REP with low hydrogen price, and (iii) a market-bidding, 1000 MW electrolyzer REP with high hydrogen price. These three cases are solved for each transmission network model, resulting in a total of nine cases. For each case, we compute the system load, cost of generation per load, system emissions per load, and total renewable power curtailment. The system-focused results are reported in Table III. We further report REP-focused results including the average electrolyzer load, REP’s profit, and the average LMP at the REP’s node in Table IV. All results are computed in the same manner as described in Section IV-D.
IV-F1 System-focused results
We first examine the system load (first column of Table III). For both hydrogen prices, the increase in system load from the base system is lower under a nodal model than under a zonal or copper-plate model. In the case of the higher hydrogen price, the consumption of the REP increases the system load by 23% under a zonal model and by 17% under a nodal model. This is expected, as our analysis in Section IV-E showed that the 1000 MW electrolyzer utilization is limited by system congestion and increased LMPs under the nodal model. However, there is no substantial further increase in system load when going from a zonal model to a copper-plate model at any hydrogen price. This suggests that congestion is primarily intra-regional and can only be fully captured by a nodal model.
@@ IV-F1 System-focused results
We conclude that the choice of transmission network model significantly affects the perceived results of adding an electrolyzer to a power system, as the added load could induce congestion in the system. Further, we highlight the role of nodal pricing in mitigating network congestion, as system congestion is captured in this price formation process, as well as the broader benefits of flexible electrolyzer operation. Our results indicate that when the electrolyzer operates flexibly and the REP actively participates in the electricity market, system congestion is either slightly reduced or does not increase significantly. This is reflected in the relatively stable total system cost of generation per unit of load for the market-bidding REPs, as shown in the rightmost plot of Fig. 7.
@@
IV-F2 REP-focused results
The results for electrolyzer utilization (first column of Table IV) further illustrate to what extent the relaxed grid models overestimate the power consumption of the REP. For example, in the case with a higher hydrogen price, the electrolyzer utilization is at 100% under the relaxed network models, compared to 73% under the nodal model. This increase in utilization contributes to an overestimation of the REP’s profits, which are reported in the second column. Further, the increase in perceived profit is in part due to the LMPs at the REP’s node (last column of Table III). The copper-plate and zonal models yield lower average LMPs, as reduced congestion allows for the dispatch of cheaper generators. In the case of the higher hydrogen price, this leads to vastly differing results on REP profitability. Specifically, the zonal and copper-plate models result in more than twice the REP’s profit compared to the nodal model. The results presented for a nodal model in Section IV-E show that a 1000 MW electrolyzer would actually yield a lower profit than a 500 MW electrolyzer in the case of a high hydrogen price, due to the increase in LMPs. This could dis-incentivize larger electrolyzer capacity investments under a nodal pricing scheme. However, the results of profits in Table IV show that this incentive is lost under the zonal and copper-plate models.
@@ IV-F2 REP-focused results
V Conclusions
This paper introduces the derivation of the opportunity-cost curve of REPs, with the aim of facilitating the integration of such plants in electricity market-clearing models and the analysis of their system-wide impacts on emissions, costs, and congestion, under perfect competition. The opportunity cost is derived from the revenue of hydrogen production and forms a curve that can be used to model REP bids in an electricity market. We compare a market-bidding REP, bidding its derived opportunity-cost curve, to a REP with a fixed electrolyzer consumption. The results are obtained by solving a DC OPF of the RTS-GMLC test system, modifying a pre-existing wind farm cost curve to represent the REP.
The results show that active market participation can efficiently mitigate increases in the cost of generation per load that occur when the electrolyzer is operated in a fixed manner. A market-bidding REP also efficiently reduces the curtailed renewable power in the system compared to the fixed electrolyzer case. However, high hydrogen prices will move the REP further up on the merit-order curve. This means that the REP’s electrolyzer will consume power at higher cleared electricity prices, which causes it to behave more like a fixed load in the system. Therefore, higher hydrogen prices reduces the system-level benefits of a market-bidding REP. Further, market-bidding in itself is not necessarily an efficient measure to prevent increases in system emissions due to the lack of correlation between the emission intensities and the marginal cost of generators.